OpenStudy (anonymous):

If a triangle has side lengths 16, 11, and 10, what kind of triangle is it and how do you find that out?

OpenStudy (anonymous):

what you are hoping for is that it is a right triangle. Use Pythagoras formula. \[c^2 = a^2+ b^2\] where c is the longest side. IF you get a true statement out of this then your triangle is a right triangle. \[16^2 = 10^2 + 11 ^2\]? \[256 = 100 + 121\]? \[256 = 221\]? no, 256 does not equal 221, so in this case it is not a right triangle

OpenStudy (anonymous):

if \[c^2 < a^2 + b^2\] then you have an acute triangle, meaning that all the angles are smaller than 90 degrees. if \[c^2 > a^2 + b^2\] then you have an obtuse triangle, meaning that 1 angle of the triangle is greater than 90 degrees. in this case, \[c^2 > a^2 + b^2\] so you have an obtuse triangle

OpenStudy (anonymous):

Thank you! c:

OpenStudy (anonymous):

anytime :)

OpenStudy (anonymous):

Obtuse scalene triangle. See picture and other properties of this triangle: http://www.triangle-calculator.com/?what=sss&a=16&b=11&c=10&submit=Solve